OpenStudy (anonymous):

A public library has fiction and nonfiction books in the ratio 3:2 If the library has 360 fiction books, how many nonfiction books does it have?

OpenStudy (anonymous):

i need help and i need to know how you got it please

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

I do it like this... 3/2 = 360/x (3 fiction to 2 non fiction = 360 fiction to x non fiction) Now cross multiply (3)(x) = (360)(2) 3x = 720 x = 720/3 x = 240 now we can check it... 3/2 = 360/240 cross multiply (3)(240) = (2)(360) 720 = 720 (correct) ANSWER : non fiction books (x) = 240